OpenStudy (anonymous):

Stuck, please help! A piece of metal has a charge of +2,5 nC. How many electrones would you have to add to neutralise the charge? I figured since one electron has a charge of 1,602*10^-19 C, the answer should be (1,602*10^-19) / (2,5*10^-9) = 6,4*10^-11 However, the correct answer is 1,6*10^10

OpenStudy (maheshmeghwal9):

u have mistaken there. u have to do reverse. just like this: - \[\frac{2.5 \times 10^{-9}}{1.602 \times 10^{-19}}.\]

OpenStudy (maheshmeghwal9):

= \[1.6 \times 10^{10}.\]

OpenStudy (maheshmeghwal9):

Did u gt @math0101 ?

OpenStudy (anonymous):

Ah, I see. That *should* be obvious. Thanks. :)

OpenStudy (maheshmeghwal9):

:)